OpenStudy (anonymous):

HELLLPP!! Which of these has the greatest area? Use for . rhombus with side lengths: 9 inches, height: 4 inches triangle with base length: inches, height: 4 inches circle with diameter: 4 inches trapezoid with base lengths: 3 inches and inches, height: 6 inches

OpenStudy (reemii):

You have to compute all the areas.. do you know the formulas?

OpenStudy (anonymous):

no

OpenStudy (reemii):

circle: \(\pi r^2\), \(r\) is the radius = 1/2 diameter triangle: side x height / 2 trapezoid: (BASIS + basis) x height / 2 rhombus: let me see

OpenStudy (reemii):

rhombus: side lenght x height.

OpenStudy (anonymous):

Okay so now what??

OpenStudy (reemii):

compute it: circle: \(\pi r^2 = \pi (4/2)^2 = 4\pi \simeq 4\times 3.14 = 12.56\) triangle: oh a value is missing in your post trapezoid: a value is missing rhombus: \(9\times 4 = 36\).

OpenStudy (reemii):

after finding the missing values and computing the areas, simply find which one is the greatest area.

OpenStudy (anonymous):

I'm so lost

OpenStudy (reemii):

do you see that in the question you posted on this page, there are numbers missing? for example for the triangle, we need two values (the basis, the height) in order to compute the height. but the basis' length is missing.